OpenStudy (anonymous):

What can the lengths of the bases of a trapezoide be with altitude of 8 cm and a area of 56 square cm? A.40cm and 72cm B.5 cm and 9 cm C.10cm and 18 cm D.14cm and 17 cm E.None

OpenStudy (anonymous):

56 = 0.5 * (a + b) * 8 a + b = 56 / 4 = 14 answer 5& 9 b

OpenStudy (turingtest):

A=(B1+B2)h/2 B1+B2=2A/h=112/8=14 B